How much force is required to accelerate a 5000 g mass at 20m/s2?
Answer by ikleyn(52786)

.
How much force is required to accelerate a 5000 g mass at 20 m/s2?
~~~~~~~~~~~~~~~~~~~~


Use the second Newton's law  F = m*a,


where F is the force in newtons; m is the mass in kilograms;  "a" is the acceleration in m/s^2.


Notice that the units must be consistent: convert 5000 grams to kilograms mass

    5000 grams = 5 kilograms.


After that, the calculation is straightforward:

    a = 5*20 = 100 newtons.    ANSWER

Solved and thoroughly explained.

-------------

For your education, see this link

https://www.physicsclassroom.com/class/newtlaws/Lesson-3/Newton-s-Second-Law


Happy learning (!)